OpenStudy (anonymous):

given the functions F(x)=1-3x^2 and g(x)=square root of 4-x. find fog

OpenStudy (rulnick):

1-3(4-x) = 1-12+3x = 3x-11

OpenStudy (rulnick):

f(g(x)) = 3x-11.

OpenStudy (rulnick):

okay?

OpenStudy (anonymous):

yes i just missed the last step